Calculate the enthalpy (∆H) for the following reaction:


1) H2O (g) + CO (g) = CO2 (g) + H2 (g)


2) C2H2 (g.) + 5 / 2O2 (g.) → 2CO2 (g.) + H2O (l.)


1
Expert's answer
2021-01-18T01:30:26-0500

Species Delta H C2H2 (g )= +226.7 CO2 (g )= -393.5 H20 (g) = -241.6

CO(g) = -99 kJ/mol.

1. ∆H= -393.5 -(-99 + (-241.6))= -52,9 kJ/mol.

2. ∆H= 2 x -393.5 + (-241.6) - 226.7 = 1 255,3 kJ/mol.